find a number that is twice as far from 40 as it is from 73
Found 3 solutions by LinnW, ewatrrr, Amilasor:
Answer by LinnW(1048)

2(40-x)=73-x
80 -2x = 73 - x
add 2x to each side
80 = 73 -x + 2x
80 = 73 +x
Subtract 73 from each side
80 - 73 = 73 + x -73
7 = x
Notice
2( 40 - 7 ) = 73 - 7
2(33) = 73 -7
66 = 66

Answer by ewatrrr(24785)

 
Hi,
find a number that is twice as far from 40 as it is from 73
40___________________x_____________73
Question states***
(x-40) = 2(73-x)
3x = 186
x = 62
CHECKING our Answer***
(62 -40) = 22 = 2(73-62)
40___________________62___________73

Answer by Amilasor(2)

2(x-73) = 40-x
2x-146=40-x
3x=186
x=62